OpenStudy (anonymous):

Can someone help me solve -4y + 17 (less than or equal to) 29,and tell me how it'd be graphed/shaded on a number line? Thanks so much in advance.

OpenStudy (anonymous):

I just tried solving it and I got y≥-3, is that right?

OpenStudy (anonymous):

\[-4y+17 \le 29\]Isolate y. \[-4y \le 12\]Giving,\[y \ge -3\]

OpenStudy (anonymous):

@eseidl ahh thank you!

OpenStudy (anonymous):